log(3) (2x - 1) = 3
Answer by stanbon(75887)

log(3) (2x - 1) = 3
---
base = 3
log = expoent = 3
result = 2x-1
---------------------
Equation:
2x-1 = 3^3
2x-1 = 27
2x = 28
x = 14
----------------
Cheers,
Stan H.